OpenStudy (anonymous):

MEDAL AND FOLLOW Trapezoid P Q R S and trapezoid W X Y Z shown below are similar. What is the perimeter of trapezoid W X Y Z? Trapezoid P Q R S and trapezoid W X Y Z are placed side-by-side. Side P Q is nine meters. Side Q R is twelve meters. Side R S is nine meters. Side S P is six meters. Side X Y is eight meters. 6 meters twenty meters twenty-four meters thirty-six meters

OpenStudy (anonymous):

@Lindsayb12

OpenStudy (anonymous):

perimeter of pqrs is 9+12+9+6 = 36 as wxyz is similar and XY = 8 and XY corresponds to QR each length of WXYZ is 2/3 of corresponding lengths in PQRS so perimeter of WXYZ = = 2/3 * 36 = 24 meters